a taxi company charges each person a flat fee of $2.50 plus an additional $.40 per quarter mile
a. write a formula to find the total cost for each fare
b. use the formula to find the cost for the 1 person to travel 12 miles

Answer by oberobic(2304)

Fixed Cost = $2.50
Variable Cost = $.40 per quarter mile * x,
where x is the distance traveled, measured in quarter miles.
y = Total Cost = FC + VC
or
y = .4x + 2.5
.
12 miles = 48 quarter miles.
y = .4*48 + 2.5
y = 19.2 + 2.5
y = 21.7 = $21.70
.
Notice that this equation looks similar to y = mx+b.
.
Done